OpenStudy (anonymous):

Find the mean,median,mode and range of each data set. 19,11,80,19,27,19,10,25,15

OpenStudy (anonymous):

Put them in order, least to greatest: 10,11,15,19,19,19,25,27,80. Then take the highest # (80) and subtract it from the lowest # (10) which gives you 70. The range of these #s is 70. For the mean add all #s together which is 225. Now divide this by 9 (for each number you added) 225/9=25. Mean=25 Mode: is the # that is listed the most, which is 19. The Median of a set of data is the middlemost number in the set. The median is also the number that is halfway into the set. To find the median, the data should first be arranged in order from least to greatest. Median: 19 Hope this helps!

OpenStudy (whpalmer4):

One addition: if you have an even number of data points, you calculate the median by taking the average of the middle two data points. For example: \[\{10,11,15,19\}\]Middle two data points are \(11,15\) so median = \(\frac{11+15}2 = 13\)